Solve the ordinary differential equation
d2y/dx2+dy/dx+y=0
1
Expert's answer
2020-03-09T12:56:51-0400

"y''+y'+y=0\\
\lambda^2+\lambda+1=0\\
D=1-4=-3=3i^2\\
\lambda_1=\frac{-1-i\sqrt3}{2}\\
\lambda_2=\frac{-1+i\sqrt3}{2}\\"

Fundamental system is

"y_1=e^{-\frac{1}{2}x}\cos\frac{\sqrt3}{2}x\\
y_2=e^{-\frac{1}{2}x}\sin\frac{\sqrt3}{2}x\\"

Solution of equation is

"y=c_1e^{-\frac{1}{2}x}\cos\frac{\sqrt3}{2}x+\\
+c_2e^{-\frac{1}{2}x}\sin\frac{\sqrt3}{2}x"